# HQ-SVC: Towards High-Quality Zero-Shot Singing Voice Conversion in Low-Resource Scenarios

### v2

@@ Introduction
Therefore, based on unified decoupled audio codec, we propose HQ-SVC (High-Quality Zero-Shot Singing Voice Conversion), a novel framework designed to achieve high-quality zero-shot singing voice conversion in low-resource scenarios. Additionally, to improve high-quality synthesis modeling, we first leverage decoupled codec’s zero-shot disentanglement capability to extract content and speaker features from the decoder’s intermediate layer outputs. We then introduce supplementary features to improve the performance of SVC. To further enhance feature disentanglement and fusion, we propose the Enhanced Voice Adaptation (EVA) module, which incorporates additional acoustic features for multi-feature fusion, introduces a speaker loss, and includes a Speaker-F0 Predictor to enhance synthesis quality. Following (ddsp-svc2023), the synthesis process is progressively optimized using Differentiable Digital Signal Processing (DDSP) (engel2020ddsp) and diffusion (ho2020ddpm) models to generate Mel spectrograms. Finally, a vocoder is employed to produce singing voices.
@@
Singing Voice Conversion (SVC) aims to transform a singer’s timbre into that of another target singer while preserving the melody and content. Early SVC methods were primarily based on Gaussian Mixture Models (GMM) (doi2012gmmsvc; kobayashi2014gmmsvc; kobayashi2015gmmsvc), which relied on parallel data. These methods require both source and target singers to perform the same song. This dependency not only demanded a large amount of parallel data but also limited the generation quality of the models. With the rise of Generative Adversarial Networks (GANs) (goodfellow2020gan), several GAN-based SVC approaches (sisman2019singan; sisman2020generative; liu2021fastsvc_cross) emerged that operate without parallel data. These methods employ adversarial training, where a generator and a discriminator engage in a game to improve the audio quality. However, GAN-based methods typically take longer to train and can suffer from issues like mode collapse, making the overall training process more challenging. SoVITS-SVC (so-vits-svc2023) builds on adversarial training by adding a flow-based model to prevent mode collapse during training. It is based on VITS (kim2021conditional_vits), which uses an end-to-end flow-based architecture with adversarial training to generate high-quality singing voices. Despite its effectiveness, it still demands substantial computational resources for training.
@@ Methods
We utilize FACodec (junaturalspeech3) as the unified disentangler. Specifically, we freeze all of the layers of the encoder and decoder of the FACodec and get the $x_{\mathrm{con}}$ and $x_{\mathrm{spk}}$ as the content and speaker features respectively. The basic architecture of FACodec follows Kumar et al. (kumar2024_improved_rvqgan) and employs the SnakeBeta activation function (lee2023bigvgan). The speaker extractor consists of several Conformer (gulati20_interspeech_conformer) blocks. For the three factorized vector quantization (FVQ) components $Q^{c}$, $Q^{p}$, $Q^{d}$, we use $N_{q_{c}}$ = 2, $N_{q_{p}}$ = 1, $N_{q_{d}}$ = 3 quantizers respectively, with all quantizers having a codebook size of 1024.
@@
As shown in Fig. 1 (b), since FACodec’s speaker features $x_{\mathrm{spk}}$ may include unrelated characteristics like style, they are mapped to speaker embeddings $e_{\mathrm{spk}}$ using an MLP. And the residual between the speaker features $x_{\mathrm{spk}}$ and speaker embeddings $e_{\mathrm{spk}}$ is computed to obtain residual style features $x_{\mathrm{sty}}$, which is expanded to get $e_{\mathrm{sty}}$. Following the approach in DDSP-SVC (ddsp-svc2023), we perform the transformation on the F0 to obtain
@@
where $e_{c}$ denotes the content embeddings, $e_{s}$ represents the style embeddings, and $f_{\alpha}(\cdot)$ and $f_{\beta}(\cdot)$ correspond to two independent linear layers. To achieve more effective fusion of features, the FiLM output is passed through a Conformer (gulati20_interspeech_conformer; vaswani2017attention) with 8-head self-attention and layer normalization, producing $e$, which is then used as input for the DDSP and diffusion models.

### v3

@@ Introduction
Therefore, based on unified decoupled audio codec, we propose HQ-SVC, a novel framework designed to achieve high-quality zero-shot singing voice conversion in low-resource scenarios. Additionally, to improve high-quality synthesis modeling, we first leverage decoupled codec’s zero-shot disentanglement capability to extract content and speaker features from the decoder’s intermediate layer outputs. We then introduce supplementary features to improve the performance of SVC. To further enhance feature disentanglement and fusion, we propose the Enhanced Voice Adaptation (EVA) module, which incorporates additional acoustic features for multi-feature fusion, introduces a speaker loss, and includes a Speaker-F0 Predictor to enhance synthesis quality. Following the approach in DDSP-SVC (ddsp-svc2023), the synthesis process is progressively optimized using Differentiable Digital Signal Processing (DDSP) (engel2020ddsp) and diffusion (ho2020ddpm) models to generate Mel spectrograms. Finally, a vocoder is employed to produce singing voices.
@@
Singing Voice Conversion (SVC) aims to transform a singer’s timbre into that of another target singer while preserving the melody and content. Early SVC methods (doi2012gmmsvc; kobayashi2014gmmsvc; kobayashi2015gmmsvc) were primarily based on Gaussian Mixture Models (GMM), which relied on parallel data. These methods require both source and target singers to perform the same song. This dependency not only demanded a large amount of parallel data but also limited the generation quality of the models. With the rise of Generative Adversarial Networks (GANs) (goodfellow2020gan), several GAN-based SVC approaches (sisman2019singan; sisman2020generative; liu2021fastsvc_cross) emerged that operate without parallel data. These methods employ adversarial training, where a generator and a discriminator engage in a game to improve the audio quality. However, GAN-based methods typically take longer to train and can suffer from issues like mode collapse, making the overall training process more challenging. SoVITS-SVC (so-vits-svc2023) builds on adversarial training by adding a flow-based model to prevent mode collapse during training. It is based on VITS (kim2021conditional_vits), which uses an end-to-end flow-based architecture with adversarial training to generate high-quality singing voices. While effective, it still requires considerable computational resources for training.
@@ Methods
We utilize FACodec (junaturalspeech3) as the unified disentangler, due to its high generation quality and strong disentanglement capability. Specifically, we freeze all of the layers of the encoder and decoder of the FACodec and get the $x_{\mathrm{con}}$ and $x_{\mathrm{spk}}$ as the content and speaker features respectively. The basic architecture of FACodec follows Kumar et al. (kumar2024_improved_rvqgan) and employs the SnakeBeta activation function (lee2023bigvgan). The speaker extractor consists of several Conformer (gulati20_interspeech_conformer) blocks. For the three factorized vector quantization (FVQ) components $Q^{c}$, $Q^{p}$, $Q^{d}$, we use $N_{q_{c}}$ = 2, $N_{q_{p}}$ = 1, $N_{q_{d}}$ = 3 quantizers respectively, with all quantizers having a codebook size of 1024.
@@
As shown in Fig. 1 (b), since FACodec’s speaker features $x_{\mathrm{spk}}$ may include unrelated characteristics like style, they are mapped to speaker embeddings $e_{\mathrm{spk}}$ using an MLP. And the residual between the speaker features $x_{\mathrm{spk}}$ and speaker embeddings $e_{\mathrm{spk}}$ is computed to obtain residual style features $x_{\mathrm{sty}}$, which is expanded to yield $e_{\mathrm{sty}}$ that matches the sequence length dimension. Following the approach in DDSP-SVC (ddsp-svc2023), we perform the transformation on the F0 to obtain
@@
where $e_{c}$ denotes the content embeddings, $e_{s}$ represents the style embeddings, and $f_{\alpha}(\cdot)$ and $f_{\beta}(\cdot)$ correspond to two independent linear layers. To achieve more effective fusion of features, the FiLM output is passed through a Conformer (gulati20_interspeech_conformer) with 8-head self-attention (vaswani2017attention) and layer normalization, producing $e$, which is then used as input for the DDSP and diffusion models.